OpenStudy (anonymous):

A sample of neon gas collected at a pressure of 531 mmHg and a temperature of 291 K has a mass of 10.2 grams. What is the volume (L) of the sample?

OpenStudy (anonymous):

I'd go with the ideal gas law here: \[PV=nRT\]You know P, R, and T. You can work out n (the number of moles of the gas).

OpenStudy (anonymous):

A more convient form of the ideal gas law in this case: \[PV = {m \over M} RT\] where m is the mass (in kg) and M is the molar mass (in kg/kmol or g/mol)

OpenStudy (anonymous):

i tried the PV=nRT I got (10.2 mol)(0.0821 l*atm)(291 k)/0.6986 atm

OpenStudy (anonymous):

is it 243.089/0.6986

OpenStudy (anonymous):

which comes out to 348.82

OpenStudy (anonymous):

521 mm-Hg is 69.461 kPa. The molecular mass of Neon is 20.18 kg/kmol. R = 8.314 kJ/kmol-K. Therefore, we can write the ideal gas law as follows. \[69.461 [kPa] *V[m^3] = {10.2 [g] \over 20.18 [g/mol]} 8.314 [{J \over mol-K}] 291 [K]\] \[V = 0.0176 [m^3]\] Since, 1 L = 10^-3 m^3. \[V = 17.6 [L]\]

OpenStudy (anonymous):

I got 17.27

OpenStudy (anonymous):

I had to convert mole to mass which is 0.5054